OpenStudy (anonymous):

3y − 2x = 11 y + 2x = 9 help!!!

OpenStudy (anonymous):

add them together : 3y + y -2x +2x =20 4y =20 because 2x-2x = 0 divide both sides by 4 to get your y value : y =5 plug your y-value into one of the equations: i'm plugging it into the second equation... y+2x = 9 5 +2x =9 -5 -5 2x = 4 divide both sides by 2 to get your x value... x = 2 so... y=5 and x=2

OpenStudy (anonymous):

tahst what i thought but i wasnt sure thank youuuuuuuuuuuuu!!!!!!!!

OpenStudy (anonymous):

your welcome...